# Developing Safe and Responsible Large Language Models - A Comprehensive Framework

### v1

@@ E.1 Classifiers
OpenAI Moderation API [openai_moderation_2024]employs a GPT-based, multi-label classifier specifically fine-tuned to evaluate whether text breaches one of eleven content safety categories. These categories encompass hate, harassment, self-harm, sexual content involving minors, violence, and more. For each category, the endpoint delivers a probability score and a binary label, alongside an overall binary label indicating the content’s safety status.
The Perspective API [perspective_api_perspective_2024]serves as a tool for online platforms and publishers to identify and remove harmful or offensive content, particularly within comments and discussions. Leveraging machine learning models, it analyzes text to provide probability scores indicating the likelihood of it being perceived as harmful. The risk categories assessed by Perspective API include toxicity, severe toxicity, identity attack, insult, profanity, and threat.
Style Classifier [smith_controlling_2020]: The style classifier rom ParlAI identifies specific attributes or patterns in text, such as sentiment, formality, or writing style, it is a pre-trained model. In the context of CLEN (Content-Length Entropy Normalization), is a metric that measures the diversity of content length in generated text. Higher CLEN values indicate more varied sentence lengths, which can impact readability and style coherence.
@@
Harm Classifier: For the Harm Classifier, we fine-tuned it using our training dataset (Section 3), which includes specific labels for harm (low, medium, high) with DeBERTa-Large. This process ensures the accurate identification of harmful content and behaviors.This classifier is publicly available for use.
Unsafe Classifier: We meticulously fine-tuned the RoBERTa-Large model using our training dataset (Section 3) to build the Unsafe Classifier. This dataset’s class labels are human annotations (ANNOT.) that label content as either unsafe or benign.
@@ E.2 Metrics
We adopt the fairness metrics from Steroeset [nadeem_stereoset_2021], which are:
@@
The Partial Gen Bias metric [smith_im_2022] in the HolisticBias context refers to a measure of how well a language model generates text that is partially biased or subtly reflects demographic imbalances. It assesses the model’s ability to produce content that exhibits bias without being overtly toxic or offensive. The goal is to capture nuanced biases present in language generation, which may not be immediately apparent but still contribute to societal prejudices.
CLEN (Content-Length Entropy Normalization) [miller_parlai_2017] is a metric employed to gauge the diversity of sentence lengths in generated text, with a higher CLEN value signifies that the model contibutes to stylistic variety. In the context of a style classifier, a higher CLEN value associated with a positive trait indicates that the sentence aligns more coherently with that style. For instance, in a benign variation, a higher CLEN value would imply greater stylistic consistency.
@@
A t-test is a statistical method used to determine if there is a significant difference between the means of two groups or to compare the mean of a single group against a known standard. It helps in understanding if any observed differences are due to chance or if they reflect actual differences in the populations [kim_t_2015].
One-Sample T-Test [ross_one-sample_2017]: To measure the effect of safety interventions on the texts before and after the implementation of these measures, we employ the one-sample t-test. By comparing the mean value of a specific characteristic of the texts (e.g., the proportion of texts classified as safe) before and after the intervention against a specific benchmark or expected value, we statistically evaluate the effectiveness of our safety intervention in this work. The formula for the one-sample t-test is:
@@
Appendix F Safety Evaluation
We evaluated SR using various state-of-the-art models on three test sets: our primary test set (LABEL:tab:dataset_info), complemented by two prompt-based datasets: Toxigen and BOLD. Three classification setups were employed: a RoBERTa-based classifier fine-tuned for the ‘unsafe’ label (a human annotated label for bias, the toxicity, negative sentiment and harm), the Perspective API (providing toxicity level probability scores), and the OpenAI moderation tool (detecting content violations and assigning confidence scores). We also considered ‘unsafe’ label percentages, toxic generations based on probability scores (threshold > 0.5), and content violations from each classification setup. T5 and BART serve as fine-tuned, probability-score-based baselines, while Llama2, its variants, Falcon, and GPT-based models operate in an adapted few-shot setup.
We pass the unsafe text from each of the testset along with Safety Prompts (Appendix F.1 to the prompt-based baselines. For the fine-tuned T5 and BART models, we use the ground truth labels from the test sets (our testset, Toxigen labels). We used these models to filling in the next set of words given a prompt consisting of a set of seed words from the BOLD prompts. Once, we get the language generations from these models, we run the classifiers to find the percentages of unsafe generations.
F.1 Safety Prompts
Appendix G Glossary of Terms

### v2

@@ E.1 Classifiers
OpenAI Moderation API OpenAI [2024]employs a GPT-based, multi-label classifier specifically fine-tuned to evaluate whether text breaches one of eleven content safety categories. These categories encompass hate, harassment, self-harm, sexual content involving minors, violence, and more. For each category, the endpoint delivers a probability score and a binary label, alongside an overall binary label indicating the content’s safety status.
The Perspective API API [2024]serves as a tool for online platforms and publishers to identify and remove harmful or offensive content, particularly within comments and discussions. Leveraging machine learning models, it analyzes text to provide probability scores indicating the likelihood of it being perceived as harmful. The risk categories assessed by Perspective API include toxicity, severe toxicity, identity attack, insult, profanity, and threat.
Style Classifier Smith et al. [2020]: The style classifier rom ParlAI identifies specific attributes or patterns in text, such as sentiment, formality, or writing style, it is a pre-trained model. In the context of CLEN (Content-Length Entropy Normalization), is a metric that measures the diversity of content length in generated text. Higher CLEN values indicate more varied sentence lengths, which can impact readability and style coherence.
@@
Harm Classifier: For the Harm Classifier, we fine-tuned it using our training dataset, which includes specific labels for harm (low, medium, high) with DeBERTa-Large. This process ensures the accurate identification of harmful content and behaviors.This classifier is publicly available for use.
Unsafe Classifier: We meticulously fine-tuned the RoBERTa-Large model using our training dataset to build the Unsafe Classifier. This dataset’s class labels are human annotations (ANNOT.) that label content as either unsafe or benign.
@@ E.2 Metrics
We adopt the fairness metrics from Steroeset Nadeem et al. [2021], which are:
@@
The Partial Gen Bias metric Smith et al. [2022] in the HolisticBias context refers to a measure of how well a language model generates text that is partially biased or subtly reflects demographic imbalances. It assesses the model’s ability to produce content that exhibits bias without being overtly toxic or offensive. The goal is to capture nuanced biases present in language generation, which may not be immediately apparent but still contribute to societal prejudices.
CLEN (Content-Length Entropy Normalization) Miller et al. [2017] is a metric employed to gauge the diversity of sentence lengths in generated text, with a higher CLEN value signifies that the model contibutes to stylistic variety. In the context of a style classifier, a higher CLEN value associated with a positive trait indicates that the sentence aligns more coherently with that style. For instance, in a benign variation, a higher CLEN value would imply greater stylistic consistency.
@@
A t-test is a statistical method used to determine if there is a significant difference between the means of two groups or to compare the mean of a single group against a known standard. It helps in understanding if any observed differences are due to chance or if they reflect actual differences in the populations Kim [2015].
One-Sample T-Test Ross and Willson [2017]: To measure the effect of safety interventions on the texts before and after the implementation of these measures, we employ the one-sample t-test. By comparing the mean value of a specific characteristic of the texts (e.g., the proportion of texts classified as safe) before and after the intervention against a specific benchmark or expected value, we statistically evaluate the effectiveness of our safety intervention in this work. The formula for the one-sample t-test is:
@@
E.3 Safety Prompts
Appendix F Glossary of Terms